OpenStudy (anonymous):

What is the solution for x in the proportion three over five equals twelve over x?

OpenStudy (anonymous):

Will fan and medal???????

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Let me explain...\[\frac{ 3 }{ 5 }=\frac{ 12 }{ x }\]Cross multiply, so: \[3 \times x=3x\]\[5\times12=60\]To undo the multiplication of 3x, we divide both sides by 3.\[3x \div 3=x\]\[60\div3=20\]Does this make sense?

OpenStudy (anonymous):

So, x=20

OpenStudy (anonymous):

Yes it does.